HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025, Page 7 of 9  Not logged in ELOG logo
New entries since:Thu Jan 1 01:00:00 1970
ID Date Author Category Subject Text Attachments
  42   Tue Apr 1 15:32:59 2025 ZacDAQNTCAP RestartedNTCAP Has been restarted (we restart every
day around 15:00 so data goes in a new folder)
  
  41   Tue Apr 1 14:54:24 2025 allGeneralNo SIS beam since 14:35    
  40   Tue Apr 1 14:25:21 2025 ruijiu AnalysisThe frequency of 194Pb82+ has been changed from 408.606MHz to 408.602MHy during last 12 hoursThe following images display the frequency
spectrum at the start of every hour from
3:00 to 14:00.
 12x 
  39   Tue Apr 1 14:23:01 2025 JGGeneralPleiades pocket position testsWe started playing with the detector pockets
of the pleiades arrays at 14:23 01.04.2025.
  
  38   Tue Apr 1 14:17:07 2025 RuijiuAnalysisPID of 194Pb82+ 2025.04.01.14.13.07datafile = "/data.local/G22-00018_00203/results/analyzers/RSA01/RSA01-2025.04.01.14.13.07.068.tiq_spectrum.npz"
filep = "/data.local/G22-00018_00203/scripts/OnlineDataAnalysisSystem/rionid_exercises/exercise_194Pb_Exp_2025/New_TEline2022-ESR-194Pb-208Pb-nostripper_524MeV.lpp"
alphap = "1.399"
 Screenshot_2025-04-01_14-16-42.png 
  37   Tue Apr 1 14:02:23 2025 Jelena B, SiddharthGeneralLassie   2025-04-01_14-01-37-464_tcl1064.png 
  36   Tue Apr 1 12:58:27 2025 Jelena B, SiddharthGeneralLassie   2025-04-01_12-57-50-275_tcl1064.png 
  35   Tue Apr 1 11:47:14 2025 Jelena B, SiddharthGeneralLassie 2025-04-01_11-46-35-517_tcl1064.png    2025-04-01_11-46-35-517_tcl1064.png 
  34   Tue Apr 1 10:34:20 2025 Jelena B, SiddharthGeneralResolved issue with SIS18Beam is back with some fluctuations.
We did not have beam from 9:34 until 10:08,
around 35 minutes. 
 2025-04-01_10-33-38-948_tcl1064.png 
  33   Tue Apr 1 10:03:56 2025 Jelena B, SiddharthGeneralTechnical issue with SIS18 From HKR info: there is a technical issue
with SIS18 and we will have no beam for next
25 minutes. 
 2025-04-01_10-02-16-454_tcl1064.png 
  32   Tue Apr 1 08:36:25 2025 Jelena B, SiddharthGeneralLassie 2025-04-01_08-37-32-155_tcl1064.png   2025-04-01_08-37-32-155_tcl1064.png 
  31   Tue Apr 1 06:53:51 2025 Jelena B,SiddharthGeneralLassie Monitor   2025-04-01_06-49-59-989_tcl1064.png 
  30   Tue Apr 1 06:31:46 2025 MichaelGeneralLassie Monitor 2025-04-01_06-31-10-211_tcl1064.png   2025-04-01_06-31-10-211_tcl1064.png 
  29   Tue Apr 1 05:58:51 2025 MichaelGeneralLassie Monitor 2025-04-01_05-58-09-122_tcl1064.png   2025-04-01_05-58-09-122_tcl1064.png 
  28   Tue Apr 1 05:29:45 2025 MichaelGeneralLassie Monitor 2025-04-01_05-29-01-845_tcl1064.png   2025-04-01_05-29-01-845_tcl1064.png 
  27   Tue Apr 1 04:59:14 2025 MichaelGeneralLassie Monitor 2025-04-01_04-58-31-351_tcl1064.png   2025-04-01_04-58-31-351_tcl1064.png 
  26   Tue Apr 1 04:27:10 2025 MichaelGeneralLassie Monitor 2025-04-01_04-25-53-171_tcl1064.png   2025-04-01_04-25-53-171_tcl1064.png 
  25   Tue Apr 1 03:55:59 2025 MichaelGeneralLassie Monitor 2025-04-01_03-55-21-183_tcl1064.png   2025-04-01_03-55-21-183_tcl1064.png 
  24   Tue Apr 1 03:52:56 2025 Shahab, Michael, Yue, Zac, CarloDAQgood filesall settings are done.
any files coming after this one is good:
  
  23   Tue Apr 1 03:50:12 2025 Shahab, Yue, Carlo, Zac, Esther, MichaelDAQTuners of the both cavities set410 MHz cavity resonance frequency tuned.
motors both on 50 mm
245 MHz cavity resonance frequency tuned.
  
ELOG V3.1.5-fc6679b